Compute the present value of $1,100 paid in three years using the following second year, and 9 percent in the third year. (Do not round Intermedlate calculations. Round your answer to 2 declmal places)
Answer #1

FV = PV(1+r)^n

1100 = PV(1+0.07)(1+0.08)(1+0.09)

1100 = PV(1.259604)

PV = 1100/1.259604

PV = 873.2903

Therefore Present Value = 873.29